University Printers has two service departments Maintenance and Personnel) and two operating departments (Printing and DeveloPlease show your work
Answer #1

Let X be Costs of Maintenance Dept and Y be Costs of Service dept.

So, by framing the algebraic equation we get,

x = 5000+(500/3000)y ...(1)

y = 12000+(1000/5000)x ...(2)

Substituing eq. 2 in 1, we get x = 5000+(5/3)(12000+(1/5)x)

Solving the above equations we get x=7241 and y=13448.

So,

Maintenance Personnel Printing developing
Costs before allocation 5000 12000 0 0
Maintenance -7241 1448 1448 4345
Personnel 2241 -13448 2241 8966
0 0 3690 13310

Please comment in case of any query regarding the solution.